OpenStudy (luigi0210):

Show step-by-step how to factor \[\LARGE x^2+12x+36\]

OpenStudy (abb0t):

Use quadratic formula if you can't factor.

OpenStudy (abb0t):

But the answer is obviously x = - 6

OpenStudy (shamil98):

there ya go

OpenStudy (shamil98):

For factoring: You need to think, what numbers ADDED make up 6 and MULTIPLIED give 36? xy = 36 x + y = 12 if you want you could solve it like that ^ using a system of equations to find the factors. but its obviously 6 so \[\Huge (x+6)(x+6) = x^2 + 12x + 36\] or simply \[\Huge (x+6)^2 \]

OpenStudy (shamil98):

added make up 12* typos...

OpenStudy (luigi0210):

Thank you ^_^